# Orthogonal-by-construction augmentation of physics-based input-output models

### v1

@@ 4.2 Comparison without orthogonal parametrization
The baseline model contains the dominant characteristics of the data-generating system; hence, for an arbitrary data sequence in $\mathcal{D}_{N}$, it holds that
@@
$\Theta_{\mathrm{b}}\times\Theta_{\mathrm{a}}=\{(\theta_{\mathrm{b}},\theta_{\mathrm{a}})\mid\phi(x_{k})\theta_{\mathrm{b}}+f^{\mathrm{ANN}}_{\theta_{\mathrm{a}}}=\\ \phi(x_{k})\theta_{\mathrm{b}}^{\ast}+\delta(x_{k}),\quad\forall x_{k}\in\mathcal{D}_{N}\}.$ (23)
@@
The error of the estimated baseline parameters in (25) is similar to the one in Donati et al. (2025), which also follows intuitively. When the dynamics of (14) are recovered on the training data, the exact baseline parameters are found only if the learning component identifies the unmodeled terms and nothing else. If the ANN learns parts that the baseline model could represent, the true baseline parameters can not be retained.
@@ 4.3 Consistency analysis
In the previous derivations, we have only assumed that the model training results in such parameters for which the exact relations of the data-generating system are recovered on the training set. Now, we will provide certain conditions under which this assumption is satisfied for $N\to\infty$, i.e., we will show consistency of the estimator following the arguments of Ljung (1978). For that, first we require that the data-generating system (1) is stable.
Condition 9 (Stable data-generating system)
@@
under any $k_{\mathrm{o}}\geq 0$, $x_{0},\tilde{x}_{0}\in\mathbb{R}^{n_{x}}$ with $\|x_{0}-\tilde{x}_{0}\|_{2}<\rho$, and $\{(u_{i},e_{i})\}_{i=0}^{\infty}\in\mathcal{W}_{[0,\infty]}$, where $\mathcal{W}_{[0,\infty]}$ denotes the $\sigma$-algebra associated with the random variables $\{(u_{i},e_{i})\}_{i=0}^{\infty}$; moreover, the random variables $y_{k}$ and $\tilde{y}_{k}$ satisfy (1) with the same $(u_{k},e_{k})$, but with $x_{k_{\mathrm{o}}}=x_{0}$ and $\tilde{x}_{k_{\mathrm{o}}}=\tilde{x}_{0}$.
@@
Condition 10 (Stable predictor)
@@
Consider the data-generating system (1) satisfying Condition 9 with a quasi-stationary $u$ independent of the white noise process $e$. Given the model set $\mathcal{M}$ defined by (9) satisfies Condition 10, then
@@
The identification criterion given by (4) satisfies Condition C1 in Ljung (1978), hence the proof of (Ljung, 1978, Lemma 3.1) applies for the considered case. $\blacksquare$
@@
Condition 12 (Persistency of excitation)
@@
As $N\to\infty$, the sample distribution of $\{e_{k}\}_{k=0}^{N-1}$ will converge to the original white noise distribution of $e_{k}$ with (finite) variance $\Sigma_{e}$. Thus, the second term in (32) is equal to zero, since $e_{k}$ is uncorrelated with $(\phi(x_{k})\theta_{\mathrm{b}}^{\ast}+\delta(x_{k})-\hat{y}_{k})$. Moreover, $N\to\infty$ also implies that $\frac{1}{N}\|E\|_{2}^{2}\to\mathrm{trace}(\Sigma_{e})$. Since the first term in (32) is non-negative,
$\lim_{N\to\infty}\frac{1}{N}\|Y-\hat{Y}\|_{2}^{2}\geq\mathrm{trace}(\Sigma_{e}),$ (33)
@@
Under the conditions of Theorem 11, Lemma 13, Conditions 4, and 12
@@
For the proof of (34), see Lemma 4.1 in Ljung (1978). Note that the applied loss function (4) fulfills Condition (4.4) in Ljung (1978). To prove (35), refer back to Theorem 6. With the conditions of Lemma 13, Assumption 5 is trivially satisfied; hence, (35) holds with Condition 4. $\blacksquare$
@@
For the noiseless case, i.e., when $e_{k}\equiv 0$, the error value for $\theta_{\mathrm{b}}$ in (21) holds for all $N$, provided that all other conditions of Theorem 14 are satisfied. In this scenario, attaining the global minimum of the cost function implies that the exact dynamics of the data-generating system are recovered on the training data. In contrast, when $e_{k}\neq 0$, this equivalence holds only in a statistical sense, which motivates the $N\to\infty$ condition in (35).
@@ 4.4 Covariance of the model parameters
Finally, we show that, due to the orthogonality between the baseline and learning components, the proposed parametrization results in zero covariance between the estimated $\theta_{\mathrm{b}}$ and $\theta_{\mathrm{a}}$ parameters. Under the conditions of the consistency results, the asymptotic distribution of $\hat{\theta}^{N}$ can be expressed w.r.t. to $\lim_{N\rightarrow\infty}\hat{\theta}^{N}=\theta^{\ast}\in\Theta^{\ast}$ as $\sqrt{N}(\hat{\theta}^{N}-\theta^{\ast})\in\mathrm{As}\ \mathcal{N}(0,P_{\theta})$ (Ljung, 1998). Then, under the considered quadratic loss function,
@@
where $\psi_{k}(\theta_{\ast})\in\mathbb{R}^{n_{y}\times n_{\theta}}$ is the Jacobian matrix [^1] Note that $\psi_{k}$ is defined with a different dimensional notation compared to Ljung (1998) for practical reasons. $\partial\hat{y}_{k}/\partial\theta$. Based on $\mathcal{D}_{N}$ and parameter estimate $\hat{\theta}^{N}$, $P_{\theta}$ can be estimated as
@@
Under the conditions of Theorem 14, there is zero covariance between the estimated baseline and learning component parameters, $\hat{\theta}_{\mathrm{b}}^{N}$ and $\hat{\theta}_{\mathrm{a}}^{N}$, respectively.
@@
where $J_{f}$ denotes the Jacobian of $\tilde{f}^{\mathrm{ANN}}_{\theta_{\mathrm{a}}}(x_{k})$ w.r.t $\theta_{\mathrm{a}}$. Since $\psi_{k}$ is computed on the training data set, $\tilde{f}^{\mathrm{ANN}}_{\theta_{\mathrm{a}}}$ is orthogonal to the subspace spanned by $\Phi$; moreover, all changes in the projected ANN output remain orthogonal to the columns of $\Phi$, hence $\phi^{\top}J_{f}=0$. Then, for all data points in $\mathcal{D}_{N}$
@@
5 Identification example
To demonstrate the capabilities of the proposed model augmentation approach, we have generated data [^2] The used data and the implementation of the method are available at: https://github.com/AIMotionLab-SZTAKI/orthogonal-IO-augm using the following NFIR system:
@@ 5 Identification example
where $\theta_{0}=0.01$, $\theta_{1}=1$, $\theta_{2}=-0.5$, $\theta_{3}=0.1$, moreover, $e_{k}~\sim\mathcal{N}(0,\sigma_{e})$. The baseline model to be augmented is described as
@@
where the baseline parameters are $\hat{\theta}_{1}$, and $\hat{\theta}_{2}$ with initial values as $\hat{\theta}_{1}^{0}=0.8$, $\hat{\theta}_{2}^{0}=0.03$. For the learning component, a simple feedforward ANN with 1 hidden layer and 16 neurons is applied, using the hyperbolic tangent (tanh) activation. The ANN parameters are initialized with the Xavier method (Glorot and Bengio, 2010).
@@
Three distinct datasets are generated for training to highlight the importance of data generation, each with $N=1024$ data points. The first approach applies a white noise input signal with $u_{k}\sim\mathcal{N}(0,0.3^{2})$ for generating 512 samples, then the rest of the data points are acquired by using the first half of the input signal multiplied by -1. This results in a symmetric training data set (denoted by $\mathcal{D}_{N}^{\mathrm{(1)}}$), since the even nonlinearities are orthogonal to odd nonlinearities, according to (17), Condition 4 is satisfied. The second set $\mathcal{D}_{N}^{(2)}$ is gathered by applying 1024 input samples directly generated by the previous distribution. This results in a dataset with, most likely, small asymmetries. However, these asymmetries will disappear, (17) will hold, for $N\to\infty$. Finally, $\mathcal{D}_{N}^{(3)}$ is acquired by using an asymmetric distribution of $u_{k}\sim\mathcal{N}(-0.01,0.3^{2})$, which will not satisfy Condition 4 with $N\to\infty$. For testing, a data set is constructed with $N_{\mathrm{test}}=1024$, using the same input distribution as in $\mathcal{D}_{N}^{(2)}$. For better demonstration of the results, the test data is kept noise-free. To minimize (4), we applied the Adam optimizer for 500 epochs, followed by the L-BFGS method for 1000 iterations, using the identification pipeline proposed in Bemporad (2025). For benchmarking, we applied the standard additive augmentation, as well as the proposed orthogonal-by-construction method, from 10 different initialization points.
@@
First, the data sets were generated using $\sigma_{e}=0$ to test the methodologies without noise. Based on the 10 different initialization points, the test errors and the estimation errors regarding $\theta_{\mathrm{b}}$ are shown in Fig. 1. Keep in mind that models were estimated on the same data during the Monte Carlo study; only the model initializations were varied. Both parametrizations with the three different training data sets have generated similar, highly accurate results near the numerical tolerance of the applied optimizer. The main advantage of the orthogonal-by-construction method is highlighted when we investigate the error of the estimated baseline parameters. For $\mathcal{D}_{N}^{(1)}$, the special persistence of excitation condition, i.e., Condition 4, is satisfied, and $\theta_{\mathrm{b}}$ converges to the physically true values (within numerical error), as shown in Fig. 2(a). When this condition is not fulfilled, the exact values of $\theta_{\mathrm{b}}$ can not be recovered, but the uniqueness of the estimated baseline parameter still holds, as shown in Fig. 2(b), and 2(c). It is worth mentioning that using different input sequences sampled from the same distribution would result in slightly different estimation results for $\mathcal{D}_{N}^{(2)}$ and $\mathcal{D}_{N}^{(3)}$. For certain ANN initializations, the standard additive structure has yielded reasonably accurate baseline parameters; on average, it provided significantly less accurate $\theta_{\mathrm{b}}$ estimates, as visible in Fig. 1. To show that the ANN with orthogonal projection learns exactly the missing quadratic function in (41) under Condition 4, Fig. 3(a) shows the outputs of the learning component for both the orthogonal parametrization and the standard additive structure (for $\mathcal{D}_{N}^{(1)}$ and $\sigma_{e}=0$). It is visible that the proposed method has recovered nearly identically the unmodeled quadratic function for all Monte Carlo runs, whereas the standard approach has generated diverse results depending on the parameter initialization; moreover, most of them exhibit distorted characteristics compared to the true unmodeled terms. We now show that the naive data generation approach used for acquiring $\mathcal{D}_{N}^{(2)}$ also fulfills Condition 4 as $N\to\infty$. This is illustrated by repeating the identification task with the orthogonal parametrization for various data lengths. The resulting baseline parameter error values are depicted in Fig. 3(b). As visible, the error converges towards zero for $\mathcal{D}_{N}^{(2)}$ as $N\to\infty$, on the other hand, it converges towards a fixed number in case of $\mathcal{D}_{N}^{(3)}$.
@@
Figure 1: Test errors and the error of the estimated baseline parameters with $\sigma_{e}=0$ for 10 Monte Carlo runs.
@@
Figure 2: Estimated baseline parameters with different model structures and training data distributions for $\sigma_{e}=0$ with 10 Monte Carlo steps. Horizontal dashed lines correspond to the physically true parameter values.
(a) $\mathcal{D}_{N}^{(1)}$, $\sigma_{e}=0$
(b) orth. param., $\sigma_{e}=0$
@@
Figure 3: Panel (a) shows the learning component outputs compared to the true unmodeled terms for 10 Monte Carlo runs. Panels (b) and (c) show the error of the estimated baseline parameters for different training data lengths. The results are averaged over 10 Monte Carlo runs for each $N$ with the bars representing the $\pm$ standard deviation.
After the noiseless scenario, the value of $\sigma_{e}$ was set to reach a signal-to-noise ratio (SNR) of 30 dB w.r.t. the measured output. As the importance of satisfying Condition 4 was demonstrated with the noiseless example, now only $\mathcal{D}_{N}^{(1)}$ is used. The results are shown in Table 1. Keep in mind that the test data set does not contain any noise for a better comparison. As expected, a slight increase is visible in the test errors compared to the noiseless scenario, but similarly to before, the orthogonal parametrization and the baseline structure provided nearly identical results considering only model accuracy. On the other hand, the orthogonal-by-construction method resulted in nearly two magnitudes more accurate baseline parameters. For this scenario, the claimed zero covariance is also validated. After training the orthogonal-by-construction structure, $\mathrm{Cov}(\hat{\theta}^{N})$ is computed using (38). The values of the asymptotic covariance matrix are illustrated in Fig. 4 with the elements that are smaller than $10^{-6}$ in absolute value, i.e., numerically zero, shown in black. The first two rows and columns correspond to the baseline parameters, hence the block matrix nature of $\hat{P}^{N}$ is visible. To showcase the results of Theorem 6, we repeated the identification task for the noisy case with different training data lengths. As shown in Fig. 3(c), the baseline parameter estimation error for the orthogonal-by-construction method clearly converges to zero as $N\to\infty$. In contrast, the standard additive model augmentation approach does not exhibit such convergence behavior, hence highlighting the consistency of the proposed parametrization regarding the baseline parameters.

### v2

@@ 4.2 Comparison with the standard parametrization
The baseline model contains the dominant characteristics of the data-generating system; hence, for the given data sequence in $\mathcal{D}_{N}$, it holds that
@@
$\Theta_{\mathrm{b}}\times\Theta_{\mathrm{a}}=\{(\theta_{\mathrm{b}},\theta_{\mathrm{a}})\mid\phi(x_{k})\theta_{\mathrm{b}}+f^{\mathrm{ANN}}_{\theta_{\mathrm{a}}}(x_{k})=\\ \phi(x_{k})\theta_{\mathrm{b}}^{\ast}+\delta(x_{k}),\quad\forall x_{k}\in\mathcal{D}_{N}\}.$ (23)
@@
The error of the estimated baseline parameters in (25) is similar to the one in donati_combining_2025, which also follows intuitively. When the dynamics of (14) are recovered on the training data, the exact baseline parameters are found only if the learning component identifies the unmodeled terms and nothing else. If the ANN learns parts that the baseline model could represent, the true baseline parameters can not be retained.
@@ 4.3 Consistency analysis
In the previous derivations, we assumed that the model training results in such parameters for which the exact relations of the data-generating system are recovered on the training set. Now, we will provide certain conditions under which this assumption is satisfied for $N\to\infty$, i.e., we will show consistency of the estimator following the arguments of ljung_convergence_1978. For that, first we require that the data-generating system (1) is stable.
Condition 10 (Stable data-generating system)
@@
under any $k_{\mathrm{o}}\geq 0$, $x_{0},\tilde{x}_{0}\in\mathbb{R}^{n_{\mathrm{x}}}$ with $\|x_{0}-\tilde{x}_{0}\|_{2}<\rho$, and $\{(u_{i},e_{i})\}_{i=0}^{\infty}\in\mathcal{W}_{[0,\infty]}$, where $\mathcal{W}_{[0,\infty]}$ denotes the $\sigma$-algebra associated with the random variables $\{(u_{i},e_{i})\}_{i=0}^{\infty}$; moreover, the random variables $y_{k}$ and $\tilde{y}_{k}$ satisfy (1) with the same $(u_{k},e_{k})$, but with $x_{k_{\mathrm{o}}}=x_{0}$ and $\tilde{x}_{k_{\mathrm{o}}}=\tilde{x}_{0}$.
@@
Condition 11 (Stable predictor)
@@
Consider the data-generating system (1) satisfying Condition 10 with a quasi-stationary $u$ independent of the white noise process $e$. Given the model set $\mathcal{M}$ defined by (9) satisfies Condition 11, then
@@
The identification criterion given by (4) satisfies Condition C1 in ljung_convergence_1978, hence the proof of (ljung_convergence_1978, Lemma 3.1) applies for the considered case. $\blacksquare$
Remark 13
In practice, Condition 11 serves only as a technical condition. After acquiring the estimate of the model parameters $\hat{\theta}$, it can be numerically checked whether the given model realization satisfies this requirement. For a strict guarantee of satisfying Condition 11, a so-called stable-by-design parametrization can be applied. See, e.g., revay_recurrent_2024; kon_unconstrained_2025. Extending these methods for the model augmentation setting is an open question and can be the topic of future research.
@@
Condition 14 (Persistency of excitation)
@@
For linear time-invariant models, the classical persistency of excitation (PE) condition that is characterized solely on the input trajectory in the open loop case, implies a fundamental informativity requirement, i.e., distinguishability of (equivalence sets of) parameters under the data (Gevers2009). For NARX-type models with linear-in-the-parameters form as (2), the PE condition requires the regressor matrix composed of nonlinear basis functions to be of full column rank. Unlike for linear systems, this condition depends not only on the input signal, but also on the resulting output trajectories. Consequently, PE requirements must be assessed at the level of the nonlinear regressor. Considering the full nonlinear model augmentation structure of (3) further complicates the PE condition. In practice, space-filling and model-based excitation designs (see, e.g., (yuhan_space-filling_2025)) can be applied to select inputs for maximizing data information content over the regressor space.
@@
As $N\to\infty$, the sample distribution of $\{e_{k}\}_{k=0}^{N-1}$ will converge to the original white noise distribution of $e_{k}$ with (finite) variance $\Sigma_{\mathrm{e}}$. Thus, the second term in (32) is equal to zero, since $e_{k}$ is uncorrelated with $(\phi(x_{k})\theta_{\mathrm{b}}^{\ast}+\delta(x_{k})-\hat{y}_{k})$. Moreover, $N\to\infty$ also implies that $\frac{1}{N}\|E\|_{2}^{2}\to\mathrm{trace}(\Sigma_{\mathrm{e}})$. Since the first term in (32) is non-negative,
$\lim_{N\to\infty}\frac{1}{N}\|Y-\hat{Y}\|_{2}^{2}\geq\mathrm{trace}(\Sigma_{\mathrm{e}}),$ (33)
@@
Under the conditions of Theorem 12 and Lemma 15 and Conditions 4 and 14,
@@
For the proof of (34), see Lemma 4.1 in ljung_convergence_1978. Note that the applied loss function (4) fulfills Condition (4.4) in ljung_convergence_1978. To prove (35), refer back to Theorem 7. With the conditions of Lemma 15, Assumption 6 is trivially satisfied; hence, (35) holds under Condition 4. $\blacksquare$
@@
For the noiseless case, i.e., when $e_{k}\equiv 0$, the error expression for $\theta_{\mathrm{b}}$ in (21) holds for all $N$ provided that all other conditions of Theorem 16 are satisfied. In this scenario, attaining the global minimum of the cost function implies that the exact dynamics of the data-generating system are recovered on the training data. In contrast, when $e_{k}\neq 0$, this equivalence holds only in a statistical sense, which motivates the asymptotic ($N\to\infty$) nature of (35).
@@ 4.4 Covariance of the model parameters
Finally, we show that, due to the orthogonality between the baseline and learning components, the proposed parametrization results in zero covariance between the estimated $\theta_{\mathrm{b}}$ and $\theta_{\mathrm{a}}$ parameters. Under the conditions of the consistency results, the asymptotic distribution of $\hat{\theta}^{N}$ can be expressed w.r.t. to $\lim_{N\rightarrow\infty}\hat{\theta}^{N}=\theta^{\ast}\in\Theta^{\ast}$ as $\sqrt{N}(\hat{\theta}^{N}-\theta^{\ast})\in\mathrm{As}\ \mathcal{N}(0,P_{\theta})$ (ljung_system_1998). Then, under the considered quadratic loss function,
@@
where $\psi_{k}(\theta_{\ast})\in\mathbb{R}^{n_{\mathrm{y}}\times n_{\theta}}$ is the Jacobian matrix [^1] Note that $\psi_{k}$ is defined with a different dimensional notation compared to ljung_system_1998 for practical reasons. $\partial\hat{y}_{k}/\partial\theta$. Based on $\mathcal{D}_{N}$ and parameter estimate $\hat{\theta}^{N}$, $P_{\theta}$ can be estimated as
@@
Under the conditions of Theorem 16 and $\Sigma_{\mathrm{e}}$ being diagonal, the covariance between the estimated baseline and learning component parameters, $\hat{\theta}_{\mathrm{b}}^{N}$ and $\hat{\theta}_{\mathrm{a}}^{N}$, is zero.
@@
where $J_{f}$ denotes the Jacobian of $\tilde{f}^{\mathrm{ANN}}_{\theta_{\mathrm{a}}}(x_{k})$ w.r.t $\theta_{\mathrm{a}}$, evaluated at $\hat{\theta}^{N}$. Since $\psi_{k}$ is computed on the training data set, $\tilde{f}^{\mathrm{ANN}}_{\theta_{\mathrm{a}}}$ is orthogonal to the subspace spanned by $\Phi$; moreover, all changes in the projected ANN output remain orthogonal to the columns of $\Phi$, hence $\phi^{\top}J_{f}=0$. Then, for all data points in $\mathcal{D}_{N}$
@@
5 Identification examples
This section presents two identification examples [^2] The used data and code for both examples are available at: https://github.com/AIMotionLab-SZTAKI/orthogonal-IO-augm that demonstrate the effectiveness of the proposed orthogonal-by-construction model augmentation structure. The first example considers a static nonlinear data-generating system and illustrates the theoretical properties derived in previous sections. The second example involves a physics-motivated simulation study, in which the proposed approach is further evaluated and compared with existing methods from the literature.
5.1 Academic simulation example
To demonstrate the capabilities of the proposed model augmentation approach, we have generated data using the following static nonlinear system:
@@ 5.1 Academic simulation example
where $\theta_{0}=0.01$, $\theta_{1}=1$, $\theta_{2}=-0.5$, $\theta_{3}=0.1$, moreover, $e_{k}~\sim\mathcal{N}(0,\sigma_{\mathrm{e}}^{2})$. We assume that, based on prior physical knowledge, a to-be-augmented baseline model is available:
@@
where the baseline parameters are $\hat{\theta}_{1}$, and $\hat{\theta}_{2}$ with initial values as $\hat{\theta}_{1}^{0}=0.8$, $\hat{\theta}_{2}^{0}=0.03$. Similar physics-based baseline models have high relevance in system identification theory (schoukens_identification_2017). For the learning component, a simple feedforward ANN with 1 hidden layer and 16 neurons is applied, using the hyperbolic tangent (tanh) activation. The ANN parameters are initialized with the Xavier method (glorot_understanding_2010).
@@
Three distinct datasets are generated for training to highlight the importance of data generation, each with $N=1024$ data points. The first approach applies a white noise input signal with $u_{k}\sim\mathcal{N}(0,0.3^{2})$ for generating 512 samples, then the rest of the data points are acquired by using the first half of the input signal multiplied by -1. This results in a symmetric training data set (denoted by $\mathcal{D}_{N}^{\mathrm{(1)}}$), since the even nonlinearities are orthogonal to odd nonlinearities, according to (17), Condition 4 is satisfied. The second set $\mathcal{D}_{N}^{(2)}$ is gathered by applying 1024 input samples directly generated by the previous distribution. This results in a dataset with, most likely, small asymmetries. However, these asymmetries will disappear, and (17) holds for $N\to\infty$. Finally, $\mathcal{D}_{N}^{(3)}$ is acquired by using an asymmetric distribution of $u_{k}\sim\mathcal{N}(-0.01,0.3^{2})$, which will not satisfy Condition 4 when $N\to\infty$. For testing, a data set is constructed with $N_{\mathrm{test}}=1024$ using the same input distribution as in $\mathcal{D}_{N}^{(2)}$. For better demonstration of the results, the test data is kept noise-free. To minimize (4), we applied the Adam optimizer for 500 epochs, followed by the L-BFGS method for 1000 iterations, using the identification pipeline proposed in bemporad_l-bfgs-b_2025. For benchmarking, we applied the standard additive augmentation, as well as the proposed orthogonal-by-construction method, using 10 different initialization points.
@@
First, the data sets were generated using $\sigma_{\mathrm{e}}^{2}=0$ to test the methodologies without noise. Based on the 10 different initialization points, the test errors and the estimation errors regarding $\theta_{\mathrm{b}}$ are shown in Fig. 1. Keep in mind that models were estimated on the same data during the Monte Carlo study; only initialization of the model parameters was varied. Both parametrizations with the three different training data sets have generated similar, highly accurate results near the numerical tolerance of the applied optimizer. The main advantage of the orthogonal-by-construction method is highlighted when we investigate the error of the estimated baseline parameters. For $\mathcal{D}_{N}^{(1)}$, the special persistence of excitation condition, i.e., Condition 4, is satisfied, and $\theta_{\mathrm{b}}$ converges to the physically true values (within numerical error), as shown in Fig. 2(a). When this condition is not fulfilled, the exact values of $\theta_{\mathrm{b}}$ can not be recovered, but the uniqueness of the estimated baseline parameter still holds, as shown in Fig. 2(b), and 2(c). It is worth mentioning that using different input sequences sampled from the same distribution would result in slightly different estimation results for $\mathcal{D}_{N}^{(2)}$ and $\mathcal{D}_{N}^{(3)}$. For certain ANN initializations, the standard additive structure has yielded reasonably accurate baseline parameters; on average, it provided significantly less accurate $\theta_{\mathrm{b}}$ estimates, as visible in Fig. 1. To show that the ANN with orthogonal projection learns exactly the missing quadratic function in (41) under Condition 4, Fig. 3(a) shows the outputs of the learning component for both the orthogonal parametrization and the standard additive structure (for $\mathcal{D}_{N}^{(1)}$ and $\sigma_{\mathrm{e}}^{2}=0$). It is visible that the proposed method has recovered nearly identically the unmodeled quadratic function for all Monte Carlo runs, whereas the standard approach has generated diverse results depending on the parameter initialization; moreover, most of them exhibit distorted characteristics compared to the true unmodeled terms. We now show that the naive data generation approach used for acquiring $\mathcal{D}_{N}^{(2)}$ also fulfills Condition 4 as $N\to\infty$. This is illustrated by repeating the identification task with the orthogonal parametrization for various data lengths. The resulting baseline parameter error values are depicted in Fig. 3(b). As visible, the error converges towards zero for $\mathcal{D}_{N}^{(2)}$ as $N\to\infty$, on the other hand, it converges towards a fixed number in case of $\mathcal{D}_{N}^{(3)}$.
@@
Figure 1: Error of the estimated baseline parameters and RMSE of the simulated model response under $\sigma_{\mathrm{e}}^{2}=0$ for 10 Monte Carlo runs.
@@
Figure 2: Estimated baseline parameters with different model structures and training data distributions for $\sigma_{\mathrm{e}}^{2}=0$ with 10 Monte Carlo steps. Horizontal dashed lines correspond to the physically true parameter values.
(a) $\mathcal{D}_{N}^{(1)}$, $\sigma_{\mathrm{e}}^{2}=0$
(b) orth. param., $\sigma_{\mathrm{e}}^{2}=0$
@@
Figure 3: Panel (a) shows the learning component outputs compared to the true unmodeled terms for 10 Monte Carlo runs. Panels (b) and (c) show the error of the estimated baseline parameters for different training data lengths. The results are averaged over 10 Monte Carlo runs for each $N$ with the bars representing the $\pm$ two times the standard deviation. These error bars are only visible for the standard augmentation approach, as the orthogonal parametrization brings uniqueness regarding the estimated baseline parameters.
To analyze estimation under noise, the value of $\sigma_{\mathrm{e}}^{2}$ was set to reach a signal-to-noise ratio (SNR) of 30 dB w.r.t. the measured output. As the importance of satisfying Condition 4 was demonstrated with the noiseless example, now only $\mathcal{D}_{N}^{(1)}$ is used. The results are shown in Table 1. Keep in mind that the test data set does not contain any noise for a better comparison, and similarly as before, only parameter initialization is varied during the Monte Carlo study, the estimation data is kept the same. As expected, a slight increase is visible in the test errors compared to the noiseless scenario, but similarly to before, the orthogonal parametrization and the baseline structure provided nearly identical results considering only model accuracy. On the other hand, the orthogonal-by-construction method resulted in nearly two magnitudes more accurate baseline parameters. For this scenario, the claimed zero covariance is also validated. After training the orthogonal-by-construction structure, $\mathrm{Cov}(\hat{\theta}^{N})$ is computed using (38). The values of the asymptotic covariance matrix are depicted in Fig. 4 with the elements that are smaller than $10^{-6}$ in absolute value, i.e., numerically zero, shown in black. The first two rows and columns correspond to the baseline parameters, hence the block matrix nature of $\hat{P}^{N}$ is visible. To showcase the results of Theorem 7, we repeated the identification task for the noisy case with different training data lengths. As shown in Fig. 3(c), the baseline parameter estimation error for the orthogonal-by-construction method clearly converges to zero as $N\to\infty$. In contrast, the standard additive model augmentation approach does not exhibit such convergence behavior, showing the consistency of the proposed parametrization regarding the baseline parameters.